# A geometric approach to informed MCMC sampling

### v1

@@ Abstract
A Riemannian geometric framework for Markov chain Monte Carlo (MCMC) is developed where using the Fisher-Rao metric on the manifold of probability density functions (pdfs) informed proposal densities for Metropolis-Hastings (MH) algorithms are constructed. We exploit the square-root representation of pdfs under which the Fisher-Rao metric boils down to the standard $L^{2}$ metric on the positive orthant of the unit hypersphere. The square-root representation allows us to easily compute the geodesic distance between densities, resulting in a straightforward implementation of the proposed geometric MCMC methodology. Unlike the random walk MH that blindly proposes a candidate state using no information about the target, the geometric MH algorithms effectively move an uninformed base density (e.g., a random walk proposal density) towards different global/local approximations of the target density. We compare the proposed geometric MH algorithm with other MCMC algorithms for various Markov chain orderings, namely the covariance, efficiency, Peskun, and spectral gap orderings. The superior performance of the geometric algorithms over other MH algorithms like the random walk Metropolis, independent MH and variants of Metropolis adjusted Langevin algorithms is demonstrated in the context of various multimodal, nonlinear and high dimensional examples. In particular, we use extensive simulation and real data applications to compare these algorithms for analyzing mixture models, logistic regression models and ultra-high dimensional Bayesian variable selection models. A publicly available R package accompanies the article.
@@ 1 Introduction
To build computationally efficient, informative MCMC algorithms that adapt to the geometry of the target, we consider a novel approach using Bhattacharyya’s (1943) ‘square-root’ representation for pdfs. Under this representation, the manifold of pdfs can be identified with the positive orthant of the unit sphere and the FR metric boils down to the standard $L^{2}$ metric (see Figure 1). This simplifies computations through the availability of explicit, closed-form expressions for useful geometric quantities like geodesic paths, distances as well as exponential and inverse-exponential maps. Thus, the proposed general-purpose, geometric MCMC algorithms, unlike MMALA and RMHMC, do not require the first and higher-order derivatives of the log target density. Also, Zanella (2020) considers point-wise informed proposals for discrete spaces, whereas here, we construct informed proposals respecting the geometry of the space of pdfs applicable to both discrete and continuous spaces.
@@
The rest of the article is organized as follows. Section 2 introduces the square-root representation and the FR Riemannian geometric framework. In Section 3, we define the informed proposal distributions constructed by ‘moving’ any base uninformed kernel in the directions of the target density and its local and global approximations. Section 4 provides some results for comparing general state space Markov chains. Then, Section 5 uses these general results to establish the superior performance of the proposed geometric MCMC algorithms over the MH algorithms based on the uninformed base kernel. Section 6 describes some methods for sampling from the geometric proposal distributions for efficient simulation using the proposed MH algorithms. In Section 7, we consider several widely used high dimensional and nonlinear models with complex, multimodal target distributions, namely mixture models, logistic models, and the Bayesian variable selection models with spike and slab priors. While the target distribution for the feature selection example is discrete, it is continuous for the other examples. The proposed geometric methods outperform several traditional and state-of-the-art MCMC algorithms in all these examples. Finally, in Section 8 we discuss possible extensions and future works. The appendix contains an alternative formulation of the geometric MCMC method and results from additional simulation studies for the Bayesian variable selection example. The paper is accompanied by an R package geommc.
@@ 3 Manifold MH proposals
Proof.
We have
$+\sin\big{(}\epsilon\theta_{i,x}[\sin(\theta_{i,x})]^{-1}\big{\|}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{\|}\big{)}\frac{\big{[}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{]}}{\big{\|}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{\|}}.$ (3)
Note that
$\quad-2\big{\langle}\sqrt{f(y|x)},\sqrt{g_{i}(y|x)}\big{\rangle}^{2}$
$=1-\big{\langle}\sqrt{f(y|x)},\sqrt{g_{i}(y|x)}\big{\rangle}^{2}.$ (4)
Using the fact that $\sin(\cos^{-1}(r))=\sqrt{1-r^{2}},$ we then have
$\epsilon\theta_{i,x}[\sin(\theta_{i,x})]^{-1}\big{\|}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{\|}=\epsilon\theta_{i,x}.$ (5)
Using (5) in (3), then we have
$\exp_{\sqrt{f}}(\epsilon\exp^{-1}_{\sqrt{f}}(\sqrt{g_{i}}))(y|x)=\cos(\epsilon\theta_{i,x})\sqrt{f(y|x)}+\sin(\epsilon\theta_{i,x})\zeta_{i}(y|x).$ (6)
Next, squaring both sides of (6) and using the fact that $2\cos(\epsilon\theta_{i,x})\sin(\epsilon\theta_{i,x})=\sin(2\epsilon\theta_{i,x})$ we have (1). Now
$\|\zeta_{i}(y|x)\|^{2}$ $=\langle\zeta_{i}(y|x),\zeta_{i}(y|x)\rangle$
$=1,$ (7)
$\zeta_{i}\in T_{\sqrt{f}}\mathcal{M}$. Thus from (3) it follows that (1) is a pdf. ∎
@@
In Section 6, we describe an efficient method for sampling from the geometric MH proposal density (8). The step size $\epsilon$ determines the amount of perturbation of the base density $f$ in the directions of $\mathcal{G}$. By changing $\epsilon$ from zero to one, the geodesic paths from $f$ to the densities in $\mathcal{G}$ are traced. Thus, smaller values of $\epsilon$ lead to higher acceptance rates. So, the single parameter $\epsilon$ provides a simple way of tuning the geometric MCMC chains based on the acceptance rate criteria. We now discuss some choices for the baseline pdf $f$ and useful global and local approximations of $\psi$. Given the current state $x$, $f(y|x)$ can be the normal/uniform density centered at the current state $x$, that is, $f$ is the proposal density often used in the RWM algorithms mentioned in the Introduction. Or, $f$ can be the proposal density used for an independent MH algorithm (Robert and Casella, 2004, chap. 7.4), in which case $f(y|x)=f(y)$. Now, we discuss possible choices for $\mathcal{G}$. For multimodal targets these can be a set of densities centered at the local modes (see Example 3 in Section 7), it can be the target density constrained on an appropriate neighborhood of the current state $x$ (see Example 5 in Section 7), these densities can be some normal approximations to the target posterior density according to the Bernstein-von Mises theorem (see Example 4 in Section 7) or some variational approximations of the target posterior density (Blei et al., 2017). Other possibilities of $g_{j}$ can be
@@
Instead of considering a single MH proposal based on a mixture of $\phi_{i,\epsilon}$ as in Algorithm 1, a mixture of MH Mtfs based on each $\phi_{i,\epsilon}$ can also be considered (see Algorithm 2 in the Appendix A). Tierney (1998) showed that Algorithm 1 dominates the later algorithm in the Peskun sense. On the other hand, per iteration computation cost of Algorithm 1 is higher than that of the later. The acceptance probability of Algorithm 1 requires computation of $\phi_{i,\epsilon}$ for all $i\in\{1,\dots,k\}$, whereas that of the mixture of MH algorithms needs computing $\phi_{i,\epsilon}$ only for the sampled $i$.
@@ 4 Ordering Markov chains
Proof.
Case $c>1$:
$v(t,P)=\frac{1}{c}v(t,P_{1})+\frac{1-c}{c}\sigma^{2}_{t}.$
Since $v(t,P_{1})\leq v(t,Q)$, we have
$v(t,P)\leq\frac{1}{c}v(t,Q)+\frac{1-c}{c}\sigma^{2}_{t}.$
Case $c\leq 1$:
$v(t,Q_{1})=\frac{1}{c}v(t,Q)+\frac{1-c}{c}\sigma^{2}_{t}.$
@@ 6 Sampling from the geometric MH proposal density
Since the proposal density $\phi_{i,\epsilon}$ in (8) is a mixture of two densities $f$ and $h_{i}$, given $\epsilon$ and $\theta_{i,x}$, sampling from (8) can be done by sampling from $f$ and $h_{i}$ with probabilities $\cos^{2}(\epsilon\theta_{i,x})$ and $\sin^{2}(\epsilon\theta_{i,x})$, respectively. Since $f$ is generally the proposal density of an uninformed MCMC algorithm already in use in practice, the ability to sample from (8) requires successfully sampling from $h_{i}$ and computing the mixing weights.
In the special case, when $\epsilon=1$, the mixing coefficients boil down to $\langle\sqrt{f},\sqrt{g_{i}}\rangle^{2}$ and $1-\langle\sqrt{f},\sqrt{g_{i}}\rangle^{2}$. Now $\theta_{i,x}$ as well as the pdf $h_{i}$ given in (9) involve $\langle\sqrt{f},\sqrt{g_{i}}\rangle$, which may be available in closed form. For example, if $f(y)=\phi_{d}(y;\mu_{1},\Sigma_{1})$ and $g(y)=\phi_{d}(y;\mu_{2},\Sigma_{2})$ where $\phi_{d}(y;\mu_{i},\Sigma_{i})$ denotes the probability density function of the $d-$ dimensional normal distribution with mean vector $\mu_{i}$, covariance matrix $\Sigma_{i}$, and evaluated at $y,i=1,2$, then
@@
where $u_{i}(y|x)=[g_{i}(y|x)+\cos^{2}(\theta_{i,x})f(y|x)]/[1+\cos^{2}(\theta_{i,x})]$ is a pdf. Thus, sampling from $h_{i}$ can be done by a rejection sampler, where a sample from $u_{i}$ is only accepted with probability $h_{i}(x)/[M_{i,x}u_{i}(x)]$ where $M_{i,x}=[1+\cos^{2}(\theta_{i,x})]/\sin^{2}(\theta_{i,x})$. Note that $M_{i,x}\geq 1$, and $M_{i,x}\approx 1$ if $\theta_{i,x}\approx\pi/2$, which is the case for early iterations of the Markov chain as the baseline pdf $f$ is generally not ‘close’ to $g_{i}$, the approximations of the target density $\psi$. On the other hand, if $\theta_{i,x}$ is small, for any choice of $\epsilon$, the mixture weight for $h_{i}$ in (8) is small. Thus, in that case, sampling from $\phi_{i,\epsilon}$ is likely done by sampling from $f$. Similarly, for small $\epsilon$, the probability weight for $h_{i}$ is low in (8). Also, for independent geometric MH algorithms, since $\theta_{i,x}$ needs to be computed only once, so is $M_{i,x}$.
@@ 7 Examples
In this section, we illustrate the performance of the proposed geometric MH algorithms in the context of five different examples. In Examples 1 and 2 the independent and random walk MH chains, respectively, are known to suffer from slow mixing, and the proposed framework is shown to turn these chains, which are not even geometrically ergodic into uniformly ergodic MCMC algorithms. Also, we use these two examples to demonstrate and compare the performance of the general rejection sampling scheme described in Section 6 with examples specific tuned rejection sampling for the geometric MH proposal density (8). Example 3 considers a bivariate multimodal target corresponding to mixtures of normals where the random walk MH schemes are known to get stuck in a local mode. As in Examples 1 and 2, the geometric MH scheme successfully turns these poor mixing random walk kernels into algorithms that efficiently move between the local modes. Example 4 considers the widely used Bayesian logistic models where the geometric MH algorithm is compared with a variety of other MH schemes, including the manifold MALA. Finally, Example 5 involves discrete target probability mass functions (pmfs) arising from a popular Bayesian variable selection model. The superiority of the geometric MH algorithm over other MH schemes is demonstrated using extensive ultra-high dimensional simulation examples as well as a real dataset from a genome-wide association study (GWAS) with close to a million markers.
@@
Thus, we can sample from $h$ by first drawing $X\sim\tilde{u}$ and then accepting it with probability $h(x)/[\tilde{M}\tilde{u}(x)]$. For sampling from $h$, the rejection sampler based on $\tilde{u}$ mentioned here is better than the general method mentioned based on the representation (6) as, in this example $M\equiv M_{i,x}=8.042>5.604=\tilde{M}$. On the other hand, for $\epsilon=0.5,\sin^{2}(\epsilon\theta)=0.0588$. Thus, for sampling from (8), $h$ is sampled less than 6% of the time. We run the geometric MH algorithm for 1000 iterations with $\epsilon=0.5$ using the general method (6) starting at two different values $-10$ and $-30$. From Figures 3 (c) and (d), we see that, unlike the independent MH, the proposed manifold MH algorithm successfully quickly moves to the modal region of the target, starting at -10 or even at -30.
Figure 3: Trace plots for one thousand iterations of the independence Metropolis chains (a) starting at -5, (b) starting at -10, and of the manifold MH chains (c) starting at -10, (d) starting at -30 for the standard normal target.
From (17) and Mengersen and Tweedie (1996), we know that the independent MH chain is not geometrically ergodic. Now, we will verify (14) for the geometric MH algorithm.
$=\frac{\sin^{2}(\epsilon\theta)}{1-\exp(-1/4)}+r(x),$
where $r(x)=\cos^{2}(\epsilon\theta)\exp(x-1/2)-[\sin^{2}(\epsilon\theta)\exp(-1/8)/\{1-\exp(-1/4)\}]\exp(x/2-1/4)$. It can be shown that $r(x)$ attains minimum at $c_{0}=2\log(\sin^{2}(\epsilon\theta)\exp(-1/8)/[2\cos^{2}(\epsilon\theta)\\ \{1-\exp(-1/4)\}])+1/2$. Thus,
which is positive for $\epsilon>0$. Thus, by Proposition 2, the geometric MH algorithm is uniformly ergodic in this example.
@@
for any $\epsilon>0$, by Proposition 2, the independent geometric MH algorithm is uniformly ergodic with $t_{2}$ baseline density. Finally, we ran the geometric MH algorithm with baseline proposal density $f=N(x,1)$ (that is, the normal RWM proposal density) for 10,000 iterations with $\epsilon=0.5$ (plot (d) in Figure 4). Unlike the other MH chains, autocorrelations of the geometric MH chains drop down to (practically) zero by five lags, revealing their fast mixing properties.
@@
Suppose the target density $\psi$ is $\psi(x)=\\ 0.5\phi_{2}(x;\mu_{1},\Sigma_{1})+0.5\phi_{2}(x;\mu_{2},\Sigma_{2})$, where $\mu_{1}=(0,0)^{\top}$, $\mu_{2}=(10,10)^{\top}$, $\Sigma_{1}=I$ and $\Sigma_{2}=2I$.
We ran the RWM chain with the normal proposal density with covariance matrix $2I$ started at $(5,5)^{\top}$ for $100,000$ iterations. The estimated acceptance rate is 42.14%. The first row in Table 1 provides the estimates of the means of the two coordinates, mean square Euclidean jump distance(MSJD), effective sample size (ESS) for the two coordinates separately, and multivariate ESS (mESS) based on this RWM chain. The MSJD for a Markov chain $\{X^{(n)}\}_{n\geq 1}$ is defined as $\sum_{i=2}^{n+1}\|X^{(i)}-X^{(i-1)}\|^{2}/n$. We use the R package mcmcse (Flegal et al., 2021) for computing ESS and mESS values. The mean estimates are far from the true value $(5,5)$. The reason is that the random chain failed to move out of the local mode at $(0,0)$ even after $100,000$ iterations. The left panel of Figure 5 shows only the first 1000 steps of the chain.
We then ran the proposed geometric random walk chain with the same baseline density as the random walk MH chain for $100,000$ iterations initialized at $(5,5)^{\top}$. In this case, we took $k=2$, $g_{1}(y)\equiv\phi_{2}(y;\mu_{1},\Sigma_{1})$ and $g_{2}(y)\equiv\phi_{2}(y;\mu_{2},\Sigma_{2})$. Thus, $\mathcal{G}=\{g_{1},g_{2}\}$ is a set of two unimodal densities centered at the different local modes. Note that since $f(y|x)=\phi_{2}(y;x,2I)$, $\langle\sqrt{f},\sqrt{g_{i}}\rangle,i=1,2$ are available in closed form. From the middle panel of Figure 5, which shows the first 1000 steps of the geometric chain, we see that the chain can move back and forth between the two modes. Table 1 shows that the geometric chain (GMC1(RW)) results in higher ESS values and much higher MSJD, demonstrating better mixing.

### v2

@@ Abstract
A Riemannian geometric framework for Markov chain Monte Carlo (MCMC) is developed where using the Fisher-Rao metric on the manifold of probability density functions (pdfs), informed proposal densities for Metropolis-Hastings (MH) algorithms are constructed. We exploit the square-root representation of pdfs under which the Fisher-Rao metric boils down to the standard $L^{2}$ metric on the positive orthant of the unit hypersphere. The square-root representation allows us to easily compute the geodesic distance between densities, resulting in a straightforward implementation of the proposed geometric MCMC methodology. Unlike the random walk MH that blindly proposes a candidate state using no information about the target, the geometric MH algorithms move an uninformed base density (e.g., a random walk proposal density) towards different global/local approximations of the target density, allowing effective exploration of the distribution simultaneously at different granular levels of the state space. We compare the proposed geometric MH algorithm with other MCMC algorithms for various Markov chain orderings, namely the covariance, efficiency, Peskun, and spectral gap orderings. The superior performance of the geometric algorithms over other MH algorithms like the random walk Metropolis, independent MH, and variants of Metropolis adjusted Langevin algorithms is demonstrated in the context of various multimodal, nonlinear, and high dimensional examples. In particular, we use extensive simulation and real data applications to compare these algorithms for analyzing mixture models, logistic regression models, spatial generalized linear mixed models and ultra-high dimensional Bayesian variable selection models. A publicly available R package accompanies the article.
@@ 1 Introduction
To build computationally efficient, informative MCMC algorithms that adapt to the geometry of the target, we consider a novel approach using Bhattacharyya’s (1943) ‘square-root’ representation for pdfs. Under this representation, the manifold of pdfs can be identified with the positive orthant of the unit sphere and the FR metric boils down to the standard $L^{2}$ metric (see Figure 1). This simplifies computations through the availability of explicit, closed-form expressions for useful geometric quantities like geodesic paths, distances as well as exponential and inverse-exponential maps. Thus, the proposed general-purpose, geometric MCMC algorithms, unlike MMALA and RMHMC, do not require the first and higher-order derivatives of the log target density. Also, Zanella (2020) considers point-wise informed proposals for discrete spaces, whereas here, we construct informed proposals that incorporate both ‘local’ moves of the base kernel and ‘global’ moves respecting the geometry of the space of pdfs applicable to both discrete and continuous spaces. The concurrent global-local steps enable moves between the modes without augmenting the state space with a ‘temperature’ variable as in a tempering scheme (Geyer, 1991; Marinari and Parisi, 1992) and simultaneous exploration of the modal regions.
@@
The rest of the article is organized as follows. Section 2 introduces the square-root representation and the FR Riemannian geometric framework. In Section 3, we define the informed proposal distributions constructed by ‘moving’ any base uninformed kernel in the directions of the target density and its local and global approximations. Section 4 provides some results for comparing general state space Markov chains. Then, Section 5 uses these general results to establish the superior performance of the proposed geometric MCMC algorithms over the MH algorithms based on the uninformed base kernel. Section 6 describes some methods for sampling from the geometric proposal distributions for efficient simulation using the proposed MH algorithms. In Section 7, we consider several widely used high dimensional and nonlinear models with complex, multimodal target distributions, namely mixture models, logistic models, spatial generalized linear mixed models (GLMMs) and the Bayesian variable selection models with spike and slab priors. While the target distribution for the feature selection example is discrete, it is continuous for the other examples. The proposed geometric methods outperform several traditional and state-of-the-art MCMC algorithms in all these examples. Finally, in Section 8 we discuss possible extensions and future works. The supplementary material contains an alternative formulation of the geometric MCMC method, proofs of the theoretical results, additional technical derivations for some MCMC algorithms, and extra plots and numerical results from additional real data analyses and simulation studies for the various examples considered in the paper.
@@ 3 Manifold MH proposals
In Section 6, we describe an efficient method for sampling from the geometric MH proposal density (3). The step size $\epsilon$ determines the amount of perturbation of the base density $f$ in the directions of $\mathcal{G}$. By changing $\epsilon$ from zero to one, the geodesic paths from $f$ to the densities in $\mathcal{G}$ are traced. Thus, smaller values of $\epsilon$ lead to higher acceptance rates. So, given the base density $f$ and the densities $g_{i}$’s, the single parameter $\epsilon$ provides a simple way of tuning the geometric MCMC chains based on the acceptance rate criteria. We now discuss some choices for the baseline pdf $f$ and useful global and local approximations of $\psi$. Given the current state $x$, $f(y|x)$ can be the normal/uniform density centered at the current state $x$, that is, $f$ is the proposal density often used in the RWM algorithms mentioned in the Introduction. Or, $f$ can be the proposal density used for an independent MH algorithm (Robert and Casella, 2004, chap. 7.4), in which case $f(y|x)=f(y)$. Now, we discuss possible choices for $\mathcal{G}$. For multimodal targets these can be a set of densities centered at the local modes (see Example 3 in Section 7), it can be the target density constrained on an appropriate neighborhood of the current state $x$ (see Example 7 in Section 7), these densities can be some normal approximations to the target posterior density (see Example 5 in Section 7) or the conditional posterior density (see Example 6 in Section 7) according to the Bernstein-von Mises theorem or some variational approximations of the target posterior density (Blei et al., 2017). Other possibilities of $g_{j}$ can be
@@
Instead of considering a single MH proposal based on a mixture of $\phi_{i,\epsilon}$ as in Algorithm 1, a mixture of MH Mtfs based on each $\phi_{i,\epsilon}$ can also be considered (see Algorithm 2 in the supplement). Tierney (1998) showed that Algorithm 1 dominates the later algorithm in the Peskun sense. On the other hand, per iteration computation cost of Algorithm 1 is higher than that of the later. The acceptance probability of Algorithm 1 requires computation of $\phi_{i,\epsilon}$ for all $i\in\{1,\dots,k\}$, whereas that of the mixture of MH algorithms needs computing $\phi_{i,\epsilon}$ only for the sampled $i$.
@@ 6 Sampling from the geometric MH proposal density
The simple mixture representation of the geometric MH proposal density $\phi_{i,\epsilon}$ in (3) implies that given $\epsilon$ and $\theta_{i,x}$, sampling from it can be done by sampling from $f$ and $h_{i}$ with probabilities $\cos^{2}(\epsilon\theta_{i,x})$ and $\sin^{2}(\epsilon\theta_{i,x})$, respectively. Since $f$ is generally the proposal density of an uninformed MCMC algorithm already in use in practice, the ability to sample from (3) requires successfully sampling from $h_{i}$ and computing the mixing weights.
In the special case, when $\epsilon=1$, the mixing coefficients boil down to $\langle\sqrt{f},\sqrt{g_{i}}\rangle^{2}$ and $1-\langle\sqrt{f},\sqrt{g_{i}}\rangle^{2}$. Now $\theta_{i,x}$ as well as the pdf $h_{i}$ given in (4) involve $\langle\sqrt{f},\sqrt{g_{i}}\rangle$, which is available in closed form if both $f$ and $g_{i}$ are normal densities. In particular, if $f(y)=\phi_{d}(y;\mu_{1},\Sigma_{1})$ and $g(y)=\phi_{d}(y;\mu_{2},\Sigma_{2})$ where $\phi_{d}(y;\mu_{i},\Sigma_{i})$ denotes the probability density function of the $d-$ dimensional normal distribution with mean vector $\mu_{i}$, covariance matrix $\Sigma_{i}$, and evaluated at $y,i=1,2$, then
@@
where $u_{i}(y|x)=[g_{i}(y|x)+\cos^{2}(\theta_{i,x})f(y|x)]/[1+\cos^{2}(\theta_{i,x})]$ is a pdf. Thus, sampling from $h_{i}$ can be done by a rejection sampler, where a sample from $u_{i}$ is only accepted with probability $h_{i}(x)/[M_{i,x}u_{i}(x)]$ where $M_{i,x}=[1+\cos^{2}(\theta_{i,x})]/\sin^{2}(\theta_{i,x})$. Also, the mixture representation of $u_{i}$ allows sampling from it by drawing from $f$ and $g_{i}$ with weights depending only on $\cos(\theta_{i,x})$. Note that $M_{i,x}\geq 1$, and $M_{i,x}\approx 1$ if $\theta_{i,x}\approx\pi/2$, which is the case for early iterations of the Markov chain as the baseline pdf $f$ is generally not ‘close’ to $g_{i}$, the approximations of the target density $\psi$. On the other hand, if $\theta_{i,x}$ is small, for any choice of $\epsilon$, the mixture weight for $h_{i}$ in (3) is small. Thus, in that case, sampling from $\phi_{i,\epsilon}$ is likely done by sampling from $f$. Similarly, for small $\epsilon$, the probability weight for $h_{i}$ is low in (3). Also, for independent geometric MH algorithms, $\theta_{i,x}$ needs to be computed only once, and so is $M_{i,x}$.
Remark 5.
The proposed rejection sampler for sampling from $h_{i}$ and hence from the geometric proposal density $\phi_{i,\epsilon}$ in (3) assumes the ability to sample from the approximate densities $g_{i}$’s and to compute the $L^{2}$ inner product $\langle\sqrt{f},\sqrt{g_{i}}\rangle,i=1,\dots,k$. In all continuous target examples considered in Section 7 we observe that the choice of Gaussian densities for $g_{i}$’s leads to effective inference using the resulting geometric MH algorithms. As mentioned before, in this case, $\langle\sqrt{f},\sqrt{g_{i}}\rangle$ is available in closed form, if $f$ is also a Gaussian density.
@@ 7 Examples
In this section, we illustrate the performance of the proposed geometric MH algorithms in the context of six different examples. In Examples 1 and 2, the independent and random walk MH chains, respectively, are known to suffer from slow mixing, and the proposed framework is shown to turn these chains, which are not even geometrically ergodic into uniformly ergodic MCMC algorithms. Also, we use these two examples to demonstrate and compare the performance of the general rejection sampling scheme described in Section 6 with examples specific tuned rejection sampling for the geometric MH proposal density (3). Example 3 considers a bivariate multimodal target corresponding to mixtures of normals where the random walk MH schemes are known to get stuck in a local mode. As in Examples 1 and 2, the geometric MH scheme successfully turns these poor mixing random walk kernels into algorithms that efficiently move between the local modes. In Example 4, a geometric MH-within-Gibbs sampler successfully finds all six modes of a non-normal, multi-modal target density where the RW-within-Gibbs algorithm stays trapped in a local mode, resulting in erroneous inferences. Example 5 considers the widely used Bayesian logistic models where the geometric MH algorithm is compared with a variety of other MH schemes, including the manifold MALA. In Example 6, these different algorithms are constructed and compared in a MH-within-Gibbs framework in the context of analyzing spatial GLMMs. Finally, Example 7 involves discrete target probability mass functions (pmfs) arising from a popular Bayesian variable selection model. The superiority of the geometric MH algorithm over other MH schemes is demonstrated using extensive ultra-high dimensional simulation examples as well as a real dataset from a genome-wide association study (GWAS) with close to a million markers.
@@
Thus, we can sample from $h$ by first drawing $X\sim\tilde{u}$ and then accepting it with probability $h(x)/[\tilde{M}\tilde{u}(x)]$. For sampling from $h$, the rejection sampler based on $\tilde{u}$ mentioned here is better than the general method mentioned based on the representation (6) as, in this example $M\equiv M_{i,x}=8.042>5.604=\tilde{M}$. On the other hand, for $\epsilon=0.5,\sin^{2}(\epsilon\theta)=0.0588$. Thus, for sampling from (3), $h$ is sampled less than 6% of the time. We run the geometric MH algorithm for 1000 iterations with $\epsilon=0.5$ using the general method (6) starting at $-30$. From Figures 3 (c), we see that, unlike the independent MH, the proposed manifold MH algorithm successfully quickly moves to the modal region of the target, starting even at -30.
Figure 3: Trace plots for one thousand iterations of the independence Metropolis chains (a) starting at -5, (b) starting at -10, and of the geometric MH chains (c) with $g=\psi$ starting at -30, and (d) $g$ as N $(0,30^{2})$ density starting at -50 for the standard normal target.
From (12) and Mengersen and Tweedie (1996), we know that the independent MH chain is not geometrically ergodic. On the other hand, uniform ergodicity of the geometric MH algorithm follows from the following Lemma.
Lemma 1.
The geometric MH chain corresponding to N (1,1) base density is uniformly ergodic.
Finally, we consider an independent geometric MH algorithm with a non-informative choice of $g$, namely, the density of N $(0,30^{2})$. Figures 3 (d) shows that like the other geometric chain, this also moves to the modal region in one step.
@@
for any $\epsilon>0$, by Proposition 2, the independent geometric MH algorithm is uniformly ergodic with $t_{2}$ baseline density. Next, we ran the geometric MH algorithm with baseline proposal density $f=N(x,1)$ (that is, the normal RWM proposal density) for 10,000 iterations with $\epsilon=0.5$ (plot (d) in Figure 4). Unlike the other MH chains, autocorrelations of the geometric MH chains drop down to (practically) zero by five lags, revealing their fast mixing properties. Finally, we consider versions of the two geometric MH algorithms used here with $g$ as N (0, 900) density. From the Figure S1 given in the supplement, we see that the geometric algorithms with this choice of $g$ also result in rapidly declining autocorrelations values.
@@
Suppose the target density $\psi$ is $\psi(x)=0.5\phi_{2}(x;\mu_{1},\Sigma_{1})+0.5\phi_{2}(x;\mu_{2},\Sigma_{2})$, where $\mu_{1}=(0,0)^{\top}$, $\mu_{2}=(10,10)^{\top}$, $\Sigma_{1}=I$ and $\Sigma_{2}=2I$.
We ran the RWM chain with the normal proposal density with covariance matrix $2I$ started at $(5,5)^{\top}$ for $100,000$ iterations. The estimated acceptance rate is 42.14%. The first row in Table 1 provides the estimates of the means of the two coordinates, mean square Euclidean jump distance(MSJD), effective sample size (ESS) for the two coordinates, the multivariate ESS (mESS) and its time normalized value (mESS/sec) based on this RWM chain. The MSJD for a Markov chain $\{X^{(n)}\}_{n\geq 1}$ is defined as $\sum_{i=2}^{n+1}\|X^{(i)}-X^{(i-1)}\|^{2}/n$. We use the R package mcmcse (Flegal et al., 2021) for computing ESS and mESS values. The mean estimates are far from the true value $(5,5)$. The reason is that the random chain failed to move out of the local mode at $(0,0)$ even after $100,000$ iterations. The left panel of Figure 5 shows only the first 1000 steps of the chain.
We then ran the proposed geometric random walk chain with the same baseline density as the random walk MH chain for $100,000$ iterations initialized at $(5,5)^{\top}$. In this case, we took $k=2$, $g_{1}(y)\equiv\phi_{2}(y;\mu_{1},\Sigma_{1})$ and $g_{2}(y)\equiv\phi_{2}(y;\mu_{2},\Sigma_{2})$. Thus, $\mathcal{G}=\{g_{1},g_{2}\}$ is a set of two unimodal densities centered at the different local modes. Note that since $f(y|x)=\phi_{2}(y;x,2I)$, $\langle\sqrt{f},\sqrt{g_{i}}\rangle,i=1,2$ are available in closed form. From the middle panel of Figure 5, which shows the first 1000 steps of the geometric chain, we see that by combining the localized steps of the RW with the global moves, the geometric MCMC chain can successfully move back and forth between the two modes and simultaneously explore the modal regions. Table 1 shows that the geometric chain (GMC1(RW)) results in higher ESS values and much higher MSJD, demonstrating better mixing and successful reduction of the RW behavior.